# Artificial Intelligence Graduate School (AIGS) Computer Science and Engineering (CSE)

### v1

@@ 1 Introduction
As an increased number of vision applications are moving towards low-end devices, a variety of on-device CNNs (convolutional neural networks) have been devised in an efficient and lightweight manner. Among many of them, MobileNet [23] has become the de facto standard for many resource-constrained embedded, mobile, and IoT devices thanks to its efficient network architecture, especially the simple yet effective bottleneck block. Although MobileNet [23] enables low-end devices to run various vision tasks, e.g., image classification [5] and object detection [9, 19], by stacking multiple bottleneck blocks in the network, its peak memory (RAM) requirement, i.e., 5.6 MB for ImageNet [5], still far exceeds the limited memory capacity of millions of low-end devices having only several hundreds of KB of RAM, e.g., 256 KB, as shown in Fig. 1. Considering that 1) embedded cameras are becoming cheaper, smaller, more powerful, and low-powered [2, 1], and 2) an explosive number of embedded cameras are being installed on low-end devices at a fast pace, an increasing number of low-end devices and cameras are expected to handle advanced vision tasks on the device.
Figure 1: The memory requirement of MobileNet [23]: its peak memory (i.e., 5.6 MB for ImageNet [5]) does not fit the limited memory of many low-end devices (e.g., 256 KB).
@@
The bottleneck block of MobileNet [23] is the key enabler for the rapid deployment of CNNs onto resource-constrained low-end devices. Based on an inverted residual structure where the skip connections are between the bottleneck layers, it efficiently learns features and representations by applying point-wise and depth-wise convolutions. However, since it deals with large-sized images with the bottleneck block that expands the intermediate features with the point-wise (expansion) convolution, its memory usage soars (5.6 MB) beyond the memory capacity of many low-end devices having only KB of memory, such as ARM core-based IoT devices [24, 25]. Therefore, to meet the growing demand for on-device vision applications running on low-end devices, the memory usage of MobileNet, composed of multiple bottleneck blocks, should be reduced substantially.
@@
In this paper, we introduce an extremely memory-efficient CNN (convolutional neural network), which dramatically reduces the memory usage of the bottleneck block [23], bringing the powerful learning capability of MobileNet [23] into a memory-constrained low-end device for various computer vision tasks, e.g., image classification [5, 3] and object detection [9, 19]. We propose three design principles that collectively construct the proposed memory-efficient (and memory-aware) CNN, of which peak memory usage can be flexibly adjusted to fit the limited memory budget of low-end devices. First, ‘input segmentation’ divides an input image into a set of patches, decreasing the memory requirement of large input images at the beginning of the network. To compensate for the possible performance degradation caused by input image segmentation, we make patches partially overlap with each other along their edges and add a central patch that takes charge of the central part of the image to learn crucial features there. Next, each segmented input patch constructs a dedicated network path, consisting of multiple bottleneck blocks, all the way through the whole network independent of each other, which we call ‘patch tunneling’, keeping the memory usage of each patch under the memory budget. At the last layer of the network, they are combined to cooperate for the target task, e.g., image classification and object detection. Lastly, ‘bottleneck reordering’ maintains the memory usage of each bottleneck to be constant by rearranging execution orders of the point-wise and depth-wise convolution operations inside the bottleneck. Unlike existing approaches that execute convolution operations layer by layer without managing the peak memory usage, it reorders them across layers in a memory-efficient manner based on the computational properties of the convolution.
@@
We implement the proposed network and deploy it onto a real low-end device, i.e., STM32H7 [25]; the code implementation is provided at an anonymous GitHub1. We conduct experiments with two classes of vision tasks, i.e., image classification (ImageNet [5] and VWW [3]) and object detection (PASCAL-VOC [9] and MS-COCO [19]). The experiment results show that the proposed network enables on-device vision tasks with extremely low memory while providing competitive model performance, e.g., 61.58% and 63.84% top-1 accuracy on ImageNet [5] using only 63 KB and 254 KB of memory, respectively, validating the efficacy of the proposed network for memory-constrained low-end devices.
@@
  • We introduce an extremely memory-efficient CNN for memory-limited low-end devices, taking only 63 KB of memory for ImageNet [5] classification, which is 89x and 3.1x lower than state-of-the-art memory-efficient CNNs, i.e., MobileNet (5.6 MB) [23] and MCUNet (196 KB) [16], respectively.
@@
  • We implement and deploy the proposed network onto a real embedded device (i.e., STM32H7 [25])111 https://anonymous.4open.science/r/ACCV_2024-3394, demonstrating that it enables extremely memory-efficient on-device vision tasks, i.e., image classification and object detection, achieving 61.58% top-1 accuracy on ImageNet [5] using only 63 KB memory.
@@ 2.1 General Techniques for Making Networks Efficient
Various methods have been proposed to improve the efficiency of neural networks. Quantization [10, 14] simplifies the precision of number formats used in a network, usually from float32 to int8, int4, int2, or even binary bit [13], intending to reduce the amount of computation. Although it can decrease memory usage along with computation, the memory reduction ratio is bounded to the number of bits quantized, usually 4x (from float32 to int8), and the model performance tends to deteriorate as more aggressive quantization is applied [18]. Pruning [27, 26] is another technique that removes unnecessary weight parameters in a network based on their importance [11]. After pruning, a network becomes streamlined, improving resource efficiency, including the memory requirement. However, it deforms the original network architecture, causing several negative impacts, such as model performance degradation, unexpected model behavior, and repeated training procedures. Even worse, the memory usage may not be reduced after pruning if the pruned weight parameters are replaced with zeros, leaving the memory usage for the related operations the same. NAS (Neural Architecture Search) [6, 12, 8] seeks the optimal network architecture, which can be applied to find a memory-efficient network by imposing a memory constraint during the search [15]. Although it automatically searches for a potentially memory-efficient network, it is not guaranteed to find the best architecture that satisfies both the model performance and memory constraint. Moreover, the search process usually takes tens or hundreds of GPU days [30], which is exorbitant in practice. Unlike those techniques orthogonal to the proposed methods, the three design principles introduced in this paper 1) construct a CNN with flexible memory reduction, not restricted by some upper bound, unlike quantization, 2) do not severely degrade the model performance, 3) readily apply to various tasks without modifying the network architecture, and 4) do not entail a long search time nor repeated training processes as NAS. Since they are orthogonal to the proposed methods, we apply one of them, i.e., quantization, to our CNNs.
2.2 MCUNet
@@
MCUNet [17, 16] is a memory-efficient CNN (convolutional neural network) proposed to perform ImageNet [5] classification under 242 KB of memory on the STM32F board [24], achieving 60.3% and 64.9% top-1 classification accuracy in their first and second implementation, respectively. The memory requirement is reduced with the in-place depth-wise convolution and patch-based inference, executed by a run-time library called TinyEngine on the device. The network architecture that fits the target memory is searched by NAS (Neural Architecture Search) [6, 12, 8] via the joint neural architecture and inference scheduling search. Although MCUNet is the first network that classifies ImageNet [5] with the tight memory budget, it relies on existing techniques such as NAS (Neural Architecture Search) [6, 12, 8] and quantization, along with a significant embedded systems engineering procedure, entailing a large amount of time, computation, and human endeavors to find the network that fits the target memory budget. However, unlike MCUNet, the proposed network does not require quantization as mandatory and still achieves a similar level of memory usage, i.e., 254 KB with the float32 format, which can be further reduced to 63 KB when applying int8 quantization. Also, given a memory budget, the proposed network can be easily and flexibly constructed for various vision tasks, e.g., image classification and object detection, by applying the proposed three design principles accordingly, without requiring significant time, computation, and engineering efforts.
@@ 3 Method: Design Principles
Figure 2: An overview of the proposed memory-efficient CNN (convolutional neural network) constructed by the three memory-aware design principles: ‘input segmentation’, ‘patch tunneling’, and ‘bottleneck reordering’.
@@ 3.1 Input Segmentation
As the size of images fed into a CNN (convolutional neural network) is usually too large to fit the limited memory of many low-end devices, we decrease the dimension of the raw input image by dividing it into several patches through the process we call ‘input segmentation’ (Fig. 3). Given an input image of dimension $[h\times w\times c]$, where $h$ is the height, $w$ is the width, and $c$ is the number of channels, the input image is segmented into $k$ patches of dimension $[\frac{h}{\sqrt{k}}\times\frac{w}{\sqrt{k}}\times c]$. For instance, the images in ImageNet [5] (i.e., $[224\times 224\times 3]$), are segmented into four patches of the dimension $[112\times 112\times 3]$ when $k=4$, reducing the memory space required for the raw image by $\frac{1}{k}=\frac{1}{4}$ by processing each patch separately.
Figure 3: The proposed ‘input segmentation’ splits the input image into $k$ patches, along with the central patch, reducing the initial memory requirement of the network into $\frac{1}{k}$.
Although such a simple segmentation can mechanically reduce the memory usage to $\frac{1}{k}$, it may degrade the model performance due to the segregation of features, making it difficult for the segmented patches to have a common view of the entire image. To compensate for such potential performance loss likely to be caused by segmentation, we add some margins along patch edges so that they can overlap slightly with each other. With the vertical and horizontal margins, denoted as $m_{h}$ and $m_{w}$, respectively, the dimension of patches then becomes:
$[({h}/{\sqrt{k}}+m_{h})\times({w}/{\sqrt{k}}+m_{w})\times c]\vspace{-4pt}$ (1)
which makes the overlapped areas of $2m_{h}$ and $2m_{w}$ between two neighboring patches over the vertical and horizontal axis, respectively. For simplicity, by assuming that an image is square, i.e., $h=w$, and $m_{h}=m_{w}$, the ratio between the original image $[h\times w\times c]$ and the patch in Eq. 1 becomes $\frac{1}{k}+\frac{2m_{h}}{h\sqrt{k}}+\frac{m_{h}^{2}}{h^{2}}$, which becomes smaller when $k$ and $h$ increase, converging to $\frac{1}{k}$ when $h\gg m_{h}$.
For further improvement, we add the $(k+1)$-th patch on top of $k$ patches, which we call the central patch, to learn critical information located at the central part of the image likely to affect the model performance significantly, resulting in a total of $k+1$ patches. It is based on the observation that objects tend to be situated at the central area of the image so it is worth having a dedicated patch for that region. The central patch overlaps entirely with other $k$ patches, learning important features in the central area at the cost of having an additional convolution kernel for it. However, since each patch is executed independently, adding the central patch can be a useful way to enhance the model performance without the peak-memory increase as shown in the experiment (Sec. 4), which can be regarded as a good trade off between the huge model performance improvement and the relatively small increase in the network size.
@@ 3.2 Patch Tunneling
The idea of ‘patch tunneling’ is that features presented at each $k+1$ individual patches divided by ‘input segmentation’ can be learned separately all the way through isolated layers of lightweight (memory-efficient) convolutions, which we call a patch tunnel, and then be effectively combined at the last layer of the network. It is different from conventional approaches that apply the convolution operation to the non-segmented entire image with a large number of convolution kernels (filters), which requires a huge memory space for storing the input and output of the convolution, instead of learning a smaller patch with a relatively small number of kernels taking a small amount of memory. With $k+1$ patches divided by ‘input segmentation’ at the first layer of the network, ‘patch tunneling’ constructs $k+1$ independent patch tunnels consisting of multiple layers of lightweight convolutions in the form of bottleneck [23], starting from each patch, such that the memory requirement of the convolution operation of each layer does not exceed the memory budget $m_{b}$ until the last layer of the network.
Given the input feature of dimension $[h\times w\times c]$ and the convolutional kernel (filter) as $[k_{h}\times k_{w}\times c_{in}\times c_{out}]$ at an arbitrary layer of a patch tunnel, the output dimension of the convolutional operation becomes $[\bigl{(}\frac{h-k_{h}+2p_{h}}{s}+1\bigl{)}\times\bigl{(}\frac{w-k_{w}+2p_{w}}{s}+1\bigl{)}\times c_{out}]$, where $s$ is the stride, and $p_{h}$ and $p_{w}$ is the vertical and horizontal padding, respectively. Since both the input and output should be stored in memory during the convolution operation, the summation of the input and output dimension, while ignoring the minute memory usage required for the convolution computation, should be maintained at most the target memory budget $m_{b}$ as:
$\underbrace{h\times w\times c}_{\text{Input}}+\underbrace{\bigl{(}\frac{h-k_{h}+2p_{h}}{s}+1\bigl{)}\times\bigl{(}\frac{w-k_{w}+2p_{w}}{s}+1\bigl{)}\times c_{out}}_{\text{Output}}\leq m_{b}\vspace{-5pt}$ (2)
which first starts from the segmented patches in Eq. 1 as the initial input $[h\times w\times c]$ at the first layer of the network and computes the corresponding convolution output that will be used as the input for the next convolution layer.
@@
By adjusting $s,k_{h}$, $w_{h}$, $p_{h}$, $p_{w}$, and $c_{out}$ accordingly, Eq. 2 becomes satisfied with $m_{b}$ for all convolution layers of each patch tunnel. Since the dimension of each patch is approximately reduced to $\frac{1}{k}$ of the original image with some margins as in Eq. 1, it is possible to find $s$, $k_{h}$, $w_{h}$, $p_{h}$, $p_{w}$, and $c_{out}$ satisfying Eq. 2 without degrading the model performance substantially. For example, the summation of the input and output in Eq. 2 becomes under $m_{b}$ = 256 KB (the size of SRAM on STM32F [24]), by setting $s=1$, $k_{h}=3$, $w_{h}=3$, $p_{h}=0$, $p_{w}=0$, and $c_{out}=3$, given the input patch of ImageNet [5] as $[\bigl{(}\frac{h}{\sqrt{k}}+m_{h}\bigl{)}\times\bigl{(}\frac{w}{\sqrt{k}}+m_{w}\bigl{)}\times c]$ with $h=224,w=224,c=3,k=4$, and $m_{h}=m_{w}=18$.

### v2

# Designing Extremely Memory-Efficient CNNs for On-device Vision Tasks

@@ 1 Introduction
As an increased number of vision applications are moving towards low-end devices, a variety of on-device CNNs have been devised in an efficient and lightweight manner. Among many of them, MobileNet [23] has become the de facto standard for many resource-constrained embedded, mobile, and IoT devices thanks to its efficient network architecture, especially the simple yet effective bottleneck block. Although MobileNet [23] enables low-end devices to run various vision tasks(e.g., image classification [5] and object detection [9, 19]), the stacked bottleneck block structure significantly contributes to its peak memory (RAM) requirement(i.e., 5.6 MB for ImageNet [5]), which is still far exceeds the limited memory capacity of millions of low-end devices having only several hundreds of KB of RAM(e.g., 256 KB), as shown in Fig. 1. Considering that 1) embedded cameras are becoming cheaper, smaller, more powerful, and low-powered [2, 1], and 2) an explosive number of embedded cameras are being installed on low-end devices at a fast pace, an increasing number of low-end devices and cameras are expected to handle advanced vision tasks on the device.
Figure 1: Peak memory usage across MobileNetV2 layers on ImageNet 224x224, reaching 5.62MB at the third. The red line indicates our network’s 256KB peak memory in fp32.
@@
The bottleneck block of MobileNet is the key enabler for the rapid deployment of CNNs onto resource-constrained low-end devices. Based on an inverted residual structure where the skip connections are between the bottleneck layers, it efficiently learns features and representations by applying point-wise and depth-wise convolutions. However, since it deals with large-sized images with the bottleneck block that expands the intermediate features with the point-wise (expansion) convolution, its memory usage soars (5.6 MB) beyond the memory capacity of many low-end devices having only KB of memory, such as ARM core-based IoT devices [24, 25]. Therefore, to meet the growing demand for on-device vision applications running on low-end devices, the memory usage of MobileNet, composed of multiple bottleneck blocks, should be reduced substantially.
@@
In this paper, we introduce an extremely memory-efficient CNN, which dramatically reduces the memory usage of the bottleneck block [23], making MobileNet’s learning capability accessible to memory-constrained low-end devices for various computer vision tasks(e.g., image classification [5, 3] and object detection [9, 19]). We propose three design principles that collectively construct the proposed memory-efficient(and memory-aware) CNN, with peak memory usage can be flexibly adjusted to fit the limited memory budget of low-end devices. First, ‘input segmentation’ divides an input image into a set of patches, reducing the memory requirement of large input images at the beginning of the network. To compensate for the possible performance degradation caused by input segmentation, we make patches partially overlap with each other along their edges and add a central patch to capture crucial features in the center of the image. Next, each segmented input patch constructs a dedicated network path, consisting of multiple bottleneck blocks, all the way through the whole network independent of each other, which we call ‘patch tunneling’, keeping the memory usage of each patch under the memory budget. At the last layer of the network, they are combined to cooperate for the target task(e.g., image classification and object detection). Lastly, ‘bottleneck reordering’ maintains the memory usage of each bottleneck to be constant by rearranging execution orders of the point-wise and depth-wise convolution operations inside the bottleneck. Unlike existing approaches that execute convolution operations layer by layer without managing the peak memory usage, it reorders them across layers in a memory-efficient manner based on the computational properties of the convolution.
@@
We implement the proposed network and deploy it onto a real low-end device, i.e., STM32H7 [25]. We conduct experiments with two classes of vision tasks, i.e., image classification (ImageNet [5] and VWW [3]) and object detection (PASCAL-VOC [9] and MS-COCO [19]). The results show that the proposed network enables on-device vision tasks with extremely low memory while providing competitive model performance, e.g., 61.58% and 63.84% top-1 accuracy on ImageNet using only 63 KB and 254 KB of memory, respectively, validating the efficacy of the proposed network for memory-constrained low-end devices.
@@
  • We introduce an extremely memory-efficient CNN for memory-limited low-end devices, taking only 63 KB of memory for ImageNet classification, which is 89x and 3.1x lower than state-of-the-art memory-efficient CNNs, i.e., MobileNet (5.6 MB) [23] and MCUNet (196 KB) [16], respectively.
@@
  • We implement and deploy the proposed network onto a real embedded device (i.e., STM32H7 [25]), demonstrating that it enables extremely memory-efficient on-device vision tasks, i.e., image classification and object detection, achieving 61.58% top-1 accuracy on ImageNet using only 63 KB memory.
@@ 2 Related Work
General Techniques for Making Networks Efficient: Various methods have been proposed to improve the efficiency of neural networks. Quantization [10, 14] simplifies the precision of number formats used in a network, usually from float32 to int8, int4, int2, or even binary bit [13], intending to reduce the amount of computation. Although it can decrease memory usage along with computation, the memory reduction ratio is bounded to the number of bits quantized, usually 4x (from float32 to int8), and the model performance tends to deteriorate as more aggressive quantization is applied [18]. Pruning [27, 26] is another technique that removes unnecessary weight parameters in a network based on their importance [11]. After pruning, a network becomes streamlined, improving resource efficiency, including the memory requirement. However, it deforms the original network architecture, causing several negative impacts, such as model performance degradation, unexpected model behavior, and repeated training procedures. Even worse, the memory usage may not be reduced after pruning if the pruned weight parameters are replaced with zeros, leaving the memory usage for the related operations the same. NAS (Neural Architecture Search) [6, 12, 8] seeks the optimal network architecture, which can be applied to find a memory-efficient network by imposing a memory constraint during the search [15]. Although it automatically searches for a potentially memory-efficient network, it is not guaranteed to find the best architecture that satisfies both the model performance and memory constraint. Moreover, the search process usually takes tens or hundreds of GPU days [30], which is exorbitant in practice. Unlike those techniques orthogonal to the proposed methods, the three design principles introduced in this paper 1) construct a CNN with flexible memory reduction, not restricted by some upper bound, unlike quantization, 2) do not severely degrade the model performance, 3) readily apply to various tasks without modifying the network architecture, and 4) do not entail a long search time nor repeated training processes as NAS. Since they are orthogonal to the proposed methods, we apply one of them, i.e., quantization, to our CNNs.
@@
MCUNet [17, 16] is a memory-efficient CNN proposed to perform ImageNet [5] classification under 242 KB of memory on the STM32F board [24], achieving 60.3% and 64.9% top-1 classification accuracy in their first and second implementation, respectively. The memory requirement is reduced with the in-place depth-wise convolution and patch-based inference, executed by a run-time library called TinyEngine on the device. The network architecture that fits the target memory is searched by NAS [6, 12, 8] via the joint neural architecture and inference scheduling search. Although MCUNet is the first network that classifies ImageNet with the tight memory budget, it relies on existing techniques such as NAS and quantization, along with a significant embedded systems engineering procedure, entailing a large amount of time, computation, and human endeavors to find the network that fits the target memory budget. However, unlike MCUNet, the proposed network does not require quantization as mandatory and still achieves a similar level of memory usage, i.e., 254 KB with the float32 format, which can be further reduced to 63 KB when applying int8 quantization. Also, given a memory budget, the proposed network can be easily and flexibly constructed for various vision tasks, e.g., image classification and object detection, by applying the proposed three design principles accordingly, without requiring significant time, computation, and engineering efforts.
@@ 3 Method: Design Principles
Figure 2: An overview of the proposed memory-efficient CNN constructed by the three memory-aware design principles: ‘input segmentation’, ‘patch tunneling’, and ‘bottleneck reordering’.
Figure 3: The proposed ‘input segmentation’ splits the input image into $k$ patches, along with the central patch. Which results in a reduction of the initial memory requirement of the network into $\frac{1}{k}$.
@@ 3.1 Input Segmentation
As the size of images fed into a CNN is usually too large to fit the limited memory of many low-end devices, we decrease the dimension of the raw input image by dividing it into several patches through a process called ‘input segmentation’ (Fig. 3). Given an input image of dimension $[h\times w\times c]$, where $h$ is the height, $w$ is the width, and $c$ is the number of channels, the input image is segmented into $k$ patches of dimension $[\frac{h}{\sqrt{k}}\times\frac{w}{\sqrt{k}}\times c]$. For instance, the images in ImageNet [5] (i.e., $[224\times 224\times 3]$), are segmented into four patches of the dimension $[112\times 112\times 3]$ when $k=4$, reducing the memory space required for the raw image by $\frac{1}{k}=\frac{1}{4}$ by processing each patch separately. Although such a simple segmentation can mechanically reduce the memory usage to $\frac{1}{k}$, it may degrade the model performance due to the segregation of features, making it difficult for the segmented patches to have a common view of the entire image. To compensate for such potential performance loss likely to be caused by segmentation, we add some margins along patch edges so that they can overlap slightly with each other. With the vertical and horizontal margins, denoted as $m_{h}$ and $m_{w}$, respectively, the dimension of patches then becomes:
$[({h}/{\sqrt{k}}+m_{h})\times({w}/{\sqrt{k}}+m_{w})\times c]$ (1)
which makes the overlapped areas of $2m_{h}$ and $2m_{w}$ between two neighboring patches over the vertical and horizontal axis, respectively. For simplicity, by assuming that an image is square, i.e., $h=w$, and $m_{h}=m_{w}$, the ratio between the original image $[h\times w\times c]$ and the patch in Eq. 1 becomes $\frac{1}{k}+\frac{2m_{h}}{h\sqrt{k}}+\frac{m_{h}^{2}}{h^{2}}$, which becomes smaller when $k$ and $h$ increase, converging to $\frac{1}{k}$ when $h\gg m_{h}$. For further improvement, we add the $(k+1)$-th patch on top of $k$ patches, which we call the central patch, to learn critical information located at the central part of the image likely to affect the model performance significantly, resulting in a total of $k+1$ patches. It is based on the observation that objects tend to be situated at the central area of the image so it is worth having a dedicated patch for that region. The central patch overlaps entirely with other $k$ patches, learning important features in the central area at the cost of having an additional convolution kernel for it. However, since each patch is executed independently, adding the central patch can be a useful way to enhance the model performance without the peak-memory increase as shown in the experiment (Sec. 4), which can be regarded as a good trade off between the huge model performance improvement and the relatively small increase in the network size.
@@ 3.2 Patch Tunneling
The idea of ‘patch tunneling’ is that features presented at each $k+1$ individual patches divided by ‘input segmentation’ can be learned separately all the way through isolated layers of lightweight (memory-efficient) convolutions, which we call a patch tunnel, and then be effectively combined at the last layer of the network. It is different from conventional approaches that apply the convolution operation to the non-segmented entire image with a large number of convolution kernels (filters), which requires a huge memory space for storing the input and output of the convolution, instead of learning a smaller patch with a relatively small number of kernels taking a small amount of memory. With $k+1$ patches divided by ‘input segmentation’ at the first layer of the network, ‘patch tunneling’ constructs $k+1$ independent patch tunnels consisting of multiple layers of lightweight convolutions in the form of bottleneck [23], starting from each patch, such that the memory requirement of the convolution operation of each layer does not exceed the memory budget $m_{b}$ until the last layer of the network. Given the input feature of dimension $[h\times w\times c]$ and the convolutional kernel (filter) as $[k_{h}\times k_{w}\times c_{in}\times c_{out}]$ at an arbitrary layer of a patch tunnel, the output dimension of the convolutional operation becomes $[\bigl{(}\frac{h-k_{h}+2p_{h}}{s}+1\bigl{)}\times\bigl{(}\frac{w-k_{w}+2p_{w}}{s}+1\bigl{)}\times c_{out}]$, where $s$ is the stride, and $p_{h}$ and $p_{w}$ is the vertical and horizontal padding, respectively. Since both the input and output should be stored in memory during the convolution operation, the summation of the input and output dimension, while ignoring the minute memory usage required for the convolution computation, should be maintained at most the target memory budget $m_{b}$ as:
$\underbrace{h\times w\times c}_{\text{Input}}+\underbrace{\bigl{(}\frac{h-k_{h}+2p_{h}}{s}+1\bigl{)}\times\bigl{(}\frac{w-k_{w}+2p_{w}}{s}+1\bigl{)}\times c_{out}}_{\text{Output}}\leq m_{b}$ (2)
@@
which first starts from the segmented patches in Eq. 1 as the initial input $[h\times w\times c]$ at the first layer of the network and computes the corresponding convolution output that will be used as the input for the next convolution layer. By adjusting $s,k_{h}$, $w_{h}$, $p_{h}$, $p_{w}$, and $c_{out}$ accordingly, Eq. 2 becomes satisfied with $m_{b}$ for all convolution layers of each patch tunnel. Since the dimension of each patch is approximately reduced to $\frac{1}{k}$ of the original image with some margins as in Eq. 1, it is possible to find $s$, $k_{h}$, $w_{h}$, $p_{h}$, $p_{w}$, and $c_{out}$ satisfying Eq. 2 without degrading the model performance substantially. For example, the summation of the input and output in Eq. 2 becomes under $m_{b}$ = 256 KB (the size of SRAM on STM32F [24]), by setting $s=1$, $k_{h}=3$, $w_{h}=3$, $p_{h}=0$, $p_{w}=0$, and $c_{out}=3$, given the input patch of ImageNet [5] as $[\bigl{(}\frac{h}{\sqrt{k}}+m_{h}\bigl{)}\times\bigl{(}\frac{w}{\sqrt{k}}+m_{w}\bigl{)}\times c]$ with $h=224,w=224,c=3,k=4$, and $m_{h}=m_{w}=18$.